OpenStudy (anonymous):

Begins “Once upon a time…” Animals can act like humans Magical items or helpers Ends “…happily ever after.” Your story must also fully cover these math concepts. This is an educational fairy tale… the best kind! Create a quadratic function with two real zeros. Solve a system of non-linear equations with a graph. Solve a system of non-linear equations with a table. Compare linear, quadratic, and exponential growth functions using a table and a graph. Submit the graph and table. Identify

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

what type of problem is this???

OpenStudy (anonymous):

Seems familiar. use f(x) = (x-a)(x-b) with a and b > 0 solve sin x = cos x with a graph tabulate sin x and cos x and find x where they are equal y=mx+b vs y = ax^2 + c vs y = g exp(h x) use simple numbers for a,b,c,g, h, m

OpenStudy (anonymous):

Supposed to be writing a fairy tale. Your story should include all of the standard elements of fairy tales. but in a math form